# Spatio-temporal Early Prediction based on Multi-objective Reinforcement Learning

### v1

@@ 4.3. Finding the Optimal Set of Policies
2:     initialized $\mathbf{H}_{-1}$ for encoder. Sample a linear preference $\boldsymbol{\omega}\sim\mathcal{D}$
@@
9:        initialized go symbol for the decoder.
@@ 4.4. Finding the Hidden Preference
Inspired by (Yang et al., 2019), we employ the hidden preference $\boldsymbol{\omega}$, which can greatly influence the policy’s decisions. Given that only scalar rewards, $\varUpsilon\in\mathbb{R}^{m\times 2}$ are available, where $m$ represent the sample size, and $\varUpsilon_{ij}$ represent reward of the object $j$ in sample $i$. We propose that the preference follows a truncated multivariate Gaussian distribution $\mathcal{D}\left(\mu_{1},\mu_{2};\beta_{1}\sigma,\beta_{2}\sigma\right)$, where the non-negative parameters $\mu_{1},\mu_{2}$ are the means with $\mu_{1}+\mu_{2}=1$, $\sigma$ is the standard deviation, and $\beta$ is the weight indicating the significance of each reward.
We first normalise the rewards to $\varUpsilon_{ij}^{\prime}$ and then convert these to variation sizes $p_{ij}$, enabling uniformity and standardisation of rewards. Subsequently, we calculate the information entropy $E_{j}$, which quantifies uncertainty in the reward distribution. The weight for object $j$, $\beta_{j}$, is derived from $E_{j}$, where a lower entropy, suggesting more predictability, results in a higher weight. The weight is calculated as follows:
(11) $\varUpsilon_{ij}^{\prime}=\frac{\varUpsilon_{ij}-\min(\varUpsilon_{j})}{\max(\varUpsilon_{j})-\min(\varUpsilon_{j})},\ p_{ij}=\frac{\varUpsilon_{ij}^{\prime}}{\sum_{i=1}^{m}\varUpsilon_{ij}^{\prime}}$
$E_{j}=-\frac{\sum_{i=1}^{m}p_{ij}\ln p_{ij}}{\ln m},\ \beta_{j}=\frac{1-E_{j}}{2-\sum E_{j}}$
In our endeavor to estimate the parameters of this Gaussian distribution, we combine the strength of policy gradient and random search. These strategies collectively aim to maximize the expected return in the target task, while maintaining the network parameters constant. This delicate balance is mathematically captured as:
(12) $\mathop{\arg\max}\limits_{\mu_{1},\mu_{2}}\mathbb{E}_{\boldsymbol{\omega}\sim\mathcal{D}}\left(\mathbb{E}_{\mathbf{s}_{t},\mathbf{a}_{t}}\left(\sum_{t=0}^{T-1}\mathbf{r}_{t}\right)\right)$
@@ 5. Experiment
In this section, we mainly introduce the datasets, baselines and specific parameter settings, etc. Experiments were conducted around comparing the predicted performance of the Spatio-Temporal Early Prediction model based on Multi-Objective reinforcement learning (STEMO) against baselines, analyzing the effect of each module, and finding the hidden preference.
@@ 5.1.2. Parameters Settings
For our proposed model, we selected parameters based on preliminary experiments and prior research. Specifically, we adopted the Adam optimizer (Kingma and Ba, 2014) with an initial learning rate of 0.001, which proved to offer a stable and efficient convergence in our preliminary experiments. We set the hidden state dimension $h$ to 12 as it provided a good balance between computational complexity and model performance. Similarly, the batch size is set to 32. The number of preferences $N_{\omega}=16$, the node embedding dimension $e=4$, the parameter $\kappa=0.005$. In addition, $p=2$, $q=0.5$, $\rho=0.5$, and $T=12$ are all inspired by prior work. The source code is available at https://github.com/coco0106/MO-STEP.
@@ 5.2.1. Performance Comparison
As shown in Table 2, we vary the average used time percentage by adjusting the preference, where the average usage time percentage, defined as the sum of the optimal time instances $t^{*}_{i}$ for all $n$ nodes divided by the total time period $T$ and multiplied by 100%, represents the proportion of optimal time in the total time. This approach allows us to explore different prediction timeliness and assess the corresponding prediction accuracy, where the accuracy is measured by the error between the ground truth and the predicted values obtained by each node at their optimal time. We observe that: (1) When considering the same average used time percentage, the STEMO model exhibits higher accuracy compared to LSTM, DCRNN, ASTGCN, and Graph-WaveNet. The accuracy of the STEMO model shows a lesser decrease when the average usage time percentage experiences an equivalent reduction. This suggests that the STEMO model, which determines the optimal time based on the network’s dynamics, can adapt to diverse scenarios, striking a balance between timeliness and accuracy in predictions. In contrast, the fixed optimal time used by LSTM, DCRNN, ASTGCN, and Graph-WaveNet might not adequately account for the dynamic nature of various scenarios. (2) When the average used time percentage is the same, the STEMO model outperforms EARLIEST in terms of accuracy. This improvement can be attributed to the STEMO model’s consideration of spatial and temporal characteristics, especially in scenarios like the METR-LA dataset, where the nodes are densely distributed and exhibit strong correlations. The STEMO model leverages a Multi-Graph Convolutional Neural network (MGCN) that incorporates multiple time-steps similarity matrices and the distance matrix. This framework effectively captures and utilises the spatio-temporal correlations among the nodes. Additionally, our experiments revealed that ASTGCN exhibits poor performance on the NYPD dataset. This outcome could be attributed to missing values within the dataset and the limited representation ability of the model.
@@
As depicted in Table 3, above-average HV and lower S metrics for the STEMO model indicate its effectiveness in managing a trade-off between accuracy and timeliness in predictions - our primary objectives in this study. A higher HV, as observed in the STEMO model, signifies not only better convergence, illustrating the proximity of the solution set to the real Pareto frontier, but also greater extensiveness, indicating a broad coverage of the solution set in the objective space. Additionally, it implies better uniformity, showing an even distribution of individual solutions in the set. This high HV value affirms our assumption that the STEMO model can effectively balance prediction accuracy while maintaining the timeliness of predictions and offer a wide array of diverse solutions. On the other hand, the S metric provides a measure of the dispersion or spread of the solutions in the objective space. A lower S, as displayed by the STEMO model, suggests that the solutions are evenly distributed. This means that the STEMO model offers a variety of optimal times for predictions, and provides a robust and versatile model for various scenarios. This aligns well with our goal of creating a model that can cater to diverse situations with reliable results across a range of prediction times.
@@ 5.2.2. Ablation Study
As seen from Table 4, the full STEMO model consistently outperforms all variants across different average used time percentages in terms of MAE, RMSE, and MAPE. For instance, when the similarity matrix is removed (’w/o similarity matrix’), the performance degrades, especially at lower average used time percentages. This suggests the importance of the similarity matrix in capturing spatio-temporal information to maintain the model’s performance even at lower time percentages. The removal of the node embedding (’w/o node embedding’) and the policy network (’w/o policy’) also leads to a decrease in performance. This emphasizes the significance of the node embedding in capturing the spatial dependencies and the policy network in adaptively determining the optimal time for predictions. Overall, these findings underline the importance of each component and how they collectively contribute to the robust performance of the STEMO model in balancing prediction accuracy and timeliness.
@@ 5.2.3. Revealing hidden preferences
We modified the parameter $\rho$ to provide vectorized rewards for encoding two different objects: timeliness and accuracy. We used two different tasks (g1, g2) and used only 100 episodes to learn the hidden preferences. The derived hidden preferences are depicted in Table 5. The learned preferences of the model are concentrated on the diagonal, indicating that they are in good agreement with the actual potential preferences. For variant g1, the model shows a strong preference (0.67) for timeliness. This result is consistent with the primary goal of g1, which emphasizes making predictions as quickly as possible. g2 primarily focuses on accuracy, as evidenced by the significantly higher preference weight (0.97) for accuracy.
@@ 6. Discussion
Although the STEMO model shows promise in handling early spatio-temporal prediction tasks, several areas merit further exploration. First, the model’s performance might be restricted if it was primarily validated on a limited set of real-world datasets, which brings into question its generalizability to different scenarios or data types. Future studies could benefit from testing the model on a more diverse datasets from varying domains. Second, the model’s complexity might pose challenges in interpretation and computational efficiency. Third, future work should aim to simplify the model or increase its interpretability without sacrificing performance. Additionally, the model’s ability to learn hidden preferences requires deeper investigation.
@@ 7. Conclusion
We propose the Spatio-Temporal Early Prediction model based on Multi-Objective reinforcement learning (STEMO) model, specifically designed to address challenges in early spatio-temporal prediction. The STEMO model leverages the Multi-Graph Convolutional Neural network (MGCN) and Gated Recurrent Unit (GRU) to capture and analyze spatio-temporal correlations. Further, it utilizes the hidden state from the encoder’s output in conjunction with node embeddings to adaptively determine the optimal time for prediction. Additionally, we also explore hidden preferences within our model. Future work will focus on further refining our model, exploring its application in other domains, and addressing any limitations encountered in this study. We believe that our research contributes significantly to the ongoing discourse in the field of spatio-temporal prediction and will serve as a foundation for future advancements.
@@ Acknowledgements.
We acknowledge the support of the Australian Research Council (ARC) Centre of Excellence for Automated Decision-Making and Society (ADM+S) (CE200100005). This research/project was undertaken with the assistance of computing resources from RACE (RMIT AWS Cloud Supercomputing) (RMAS00045). Flora Salim would also like to acknowledge edge the support of Cisco’s National Industry Innovation Network (NIIN) Research Chair Program.

### v3

@@ 4.3. Finding the Optimal Set of Policies
2:     initialised $\mathbf{H}_{-1}$ for encoder. Sample a linear preference $\boldsymbol{\omega}\sim\mathcal{D}$
@@
9:        initialised go symbol for the decoder.
@@ 4.4. Finding the Hidden Preference
Inspired by (Yang et al., 2019), we employ the hidden preference $\boldsymbol{\omega}$, which can greatly influence the policy’s decisions. In the domain of Multi-Objective Reinforcement Learning (MORL), the concept of hidden preferences refers to the underlying, often undisclosed preferences that influence the prioritisation and weighting of multiple competing objectives within a decision-making process. These preferences are deemed ’hidden’ as they are not explicitly known a priori and must be inferred from the interactions within the environment. The elucidation of these preferences is pivotal for adapting and optimising policies across diverse tasks where the relative importance of objectives is not directly observable. This necessity arises from the inherent complexity and variability of real-world scenarios, where explicit preference delineation is impractical, thus requiring algorithms capable of discerning and adapting to these latent preferences effectively.
We search the hidden preferences with five steps, (1) The algorithm learns a single parametric policy network that is optimised over the entire space of preferences. This is significant because it allows the system to adapt its policy based on any given preference set, without needing to learn a new policy from scratch. (2) Employing a generalised version of the Bellman equation, which incorporates preferences into the decision-making process. This adaptation enables the system to handle multiple objectives by adjusting the policy according to a set of linear preferences, which can represent the relative importance of each objective. (3) The algorithm uses hindsight experience replay to learn from past decisions. (4) When provided with just scalar rewards and unknown preferences, the algorithm utilises a combination of policy gradient methods and stochastic search over the preference parameters. This approach helps in estimating the hidden preferences that would maximise the expected return, given the observed outcomes.
@@ 5. Experiment
In this section, we mainly introduce the datasets, baselines and specific parameter settings, etc. Experiments were conducted around comparing the predicted performance of the Early Spatio-Temporal Forecasting model based on Multi-Objective reinforcement learning (STEMO) against baselines, analyzing the effect of each module, and finding the hidden preference.
@@ 5.1.2. Parameters Settings
For our proposed model, we selected parameters based on preliminary experiments and prior research. Specifically, we adopted the Adam optimiser (Kingma and Ba, 2014) with an initial learning rate of 0.001, which proved to offer a stable and efficient convergence in our preliminary experiments. We set the hidden state dimension $h$ to 12 as it provided a good balance between computational complexity and model performance. Similarly, the batch size is set to 32. The number of preferences $N_{\omega}=16$, the node embedding dimension $e=4$, the parameter $\kappa=0.005$. In addition, $p=2$, $q=0.5$, $\rho=0.5$, and $T=12$ are all inspired by prior work. The source code is available at https://github.com/coco0106/MO-STEP.
@@ 5.2.1. Performance Comparison
As shown in Table 2, we vary the average used time percentage by adjusting the preference, where the average usage time percentage, defined as the sum of the optimal time instances $t^{*}_{i}$ for all $n$ nodes divided by the total time period $T$ and multiplied by 100%, represents the proportion of optimal time in the total time. This approach allows us to explore different prediction timeliness and assess the corresponding prediction accuracy, where the accuracy is measured by the error between the ground truth and the predicted values obtained by each node at their optimal time. We observe that: (1) When considering the same average used time percentage, the STEMO model exhibits higher accuracy compared to LSTM, DCRNN, ASTGCN, and Graph-WaveNet. The accuracy of the STEMO model shows a lesser decrease when the average usage time percentage experiences an equivalent reduction. This suggests that the STEMO model, which determines the optimal time based on the network’s dynamics, can adapt to diverse scenarios, striking a balance between timeliness and accuracy in forecasting. In contrast, the fixed optimal time used by LSTM, DCRNN, ASTGCN, and Graph-WaveNet might not adequately account for the dynamic nature of various scenarios. (2) When the average used time percentage is the same, the STEMO model outperforms EARLIEST in terms of accuracy. This improvement can be attributed to the STEMO model’s consideration of spatial and temporal characteristics, especially in scenarios like the METR-LA dataset, where the nodes are densely distributed and exhibit strong correlations. The STEMO model leverages a Multi-Graph Convolutional Neural network (MGCN) that incorporates multiple time-steps similarity matrices and the distance matrix. This framework effectively captures and utilises the spatio-temporal correlations among the nodes. Additionally, our experiments revealed that ASTGCN exhibits poor performance on the NYPD dataset. This outcome could be attributed to missing values within the dataset and the limited representation ability of the model.
@@
As depicted in Table 3, above-average HV and lower S metrics for the STEMO model indicate its effectiveness in managing a trade-off between accuracy and timeliness in predictions - our primary objectives in this study. A higher HV, as observed in the STEMO model, signifies not only better convergence, illustrating the proximity of the solution set to the real Pareto frontier, but also greater extensiveness, indicating a broad coverage of the solution set in the objective space. Additionally, it implies better uniformity, showing an even distribution of individual solutions in the set. This high HV value affirms our assumption that the STEMO model can effectively balance prediction accuracy while maintaining the timeliness of predictions and offer a wide array of diverse solutions. On the other hand, the S metric provides a measure of the dispersion or spread of the solutions in the objective space. A lower S, as displayed by the STEMO model, suggests that the solutions are evenly distributed. This means that the STEMO model offers a variety of optimal times for forecasting, and provides a robust and versatile model for various scenarios. This aligns well with our goal of creating a model that can cater to diverse situations with reliable results across a range of forecasting times.
@@ 5.2.2. Ablation Study
As seen from Table 4, the full STEMO model consistently outperforms all variants across different average used time percentages in terms of MAE, RMSE, and MAPE. For instance, when the similarity matrix is removed (’w/o similarity matrix’), the performance degrades, especially at lower average used time percentages. This suggests the importance of the similarity matrix in capturing spatio-temporal information to maintain the model’s performance even at lower time percentages. The removal of the node embedding (’w/o node embedding’) and the policy network (’w/o policy’) also leads to a decrease in performance. This emphasises the significance of the node embedding in capturing the spatial dependencies and the policy network in adaptively determining the optimal time for predictions. Overall, these findings underline the importance of each component and how they collectively contribute to the robust performance of the STEMO model in balancing prediction accuracy and timeliness.
@@ 5.2.3. Revealing hidden preferences
We modified the parameter $\rho$ to provide vectorised rewards for encoding two different objects: timeliness and accuracy. We used two different tasks (g1, g2) and used only 100 episodes to learn the hidden preferences. The derived hidden preferences are depicted in Table 5. The learned preferences of the model are concentrated on the diagonal, indicating that they are in good agreement with the actual potential preferences. For variant g1, the model shows a strong preference (0.67) for timeliness. This result is consistent with the primary goal of g1, which emphasizes making predictions as quickly as possible. g2 primarily focuses on accuracy, as evidenced by the significantly higher preference weight (0.97) for accuracy.
@@ 6. Discussion
Although the STEMO model shows promise in handling early spatio-temporal forecasting tasks, several areas merit further exploration. First, the model’s performance might be restricted if it was primarily validated on a limited set of real-world datasets, which brings into question its generalizability to different scenarios or data types. Future studies could benefit from testing the model on a more diverse datasets from varying domains. Second, the model’s complexity might pose challenges in interpretation and computational efficiency. Third, future work should aim to simplify the model or increase its interpretability without sacrificing performance. Additionally, the model’s ability to learn hidden preferences requires deeper investigation.
@@ 7. Conclusion
We propose the Early Spatio-Temporal Forecasting model based on Multi-Objective reinforcement learning (STEMO) model, specifically designed to address challenges in early spatio-temporal forecasting. The STEMO model leverages the Multi-Graph Convolutional Neural network (MGCN) and Gated Recurrent Unit (GRU) to capture and analyze spatio-temporal correlations. Further, it utilizes the hidden state from the encoder’s output in conjunction with node embeddings to adaptively determine the optimal time for forecasting. Additionally, we also explore hidden preferences within our model. Future work will focus on further refining our model, exploring its application in other domains, and addressing any limitations encountered in this study. We believe that our research contributes significantly to the ongoing discourse in the field of spatio-temporal forecasting and will serve as a foundation for future advancements.
@@ Acknowledgements.
We acknowledge the support of the Australian Research Council (ARC) Centre of Excellence for Automated Decision-Making and Society (ADM+S) (CE200100005). We also acknowledge the support of National Natural Science Foundation of China (U21A20446). This research/project was undertaken with the assistance of computing resources from RACE (RMIT AWS Cloud Supercomputing) (RMAS00045). Flora Salim would also like to acknowledge edge the support of Cisco’s National Industry Innovation Network (NIIN) Research Chair Program.